How many different ways can 4 women and 7 men be selected from
11 married couples (assuming no gay marriages).

Answer by Edwin McCravy(20054)

"11 women choose 4", that's C(11,4) or 330.

For each of those 330 ways to choose the 4 women, there are

"11 men choose 7", that's C(11,7) or 330 ways to choose the men.

Therefore the answer is C(11,4)·C(11,7) = 330·330 = 108,900 ways.

Edwin